The aluminum in a package containing 75 ft² of kitchen foil weighs approximately 12 ounces. Aluminum has a density of 2.70 g/cm³. What is the approximate thickness of the aluminum foil in millimeters? (1 ounce = 28.4 g)


1
Expert's answer
2021-04-05T07:51:30-0400

First: 12 oz X 28.4 g/oz = 340.8 g

Second: 340.8 g X 1cm3/2.7g = 126.22 cm3

Third: 75 ft2 X 12 in/ft X 12 in/ft X 2.54 cm/in X 2.54 cm/in = 69,677.28 cm2

Fourth: 69,677.28 cm2 X Thickness = 126.22 cm3

Solve for Thickness: 126.22 cm3 / 69,677.28 cm2 = 0.00181 cm

Last: 0.00181 cm X 10 mm/1 cm = 0.0181 mm